40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 29 02:17:40 2012, yuta, Update, IOO, prep for the vent - beam attenuating 
    Reply  Fri Jun 29 03:21:32 2012, yuta, Update, General, prep for the vent - others MEDMscreenshotswithCOW_20120629.png
       Reply  Fri Jun 29 11:02:00 2012, steve, Update, VAC, the vent is at 500 torr sussum.pngpzt.png
          Reply  Fri Jun 29 13:11:31 2012, steve, Update, VAC, 40m IFO at atm vent06292012.png40m@ATM.png
Message ID: 6893     Entry time: Fri Jun 29 03:21:32 2012     In reply to: 6892     Reply to this: 6894
Author: yuta 
Type: Update 
Category: General 
Subject: prep for the vent - others 

1. Turned off high voltage power supplies for PZT1/2 (input PZTs) and OMC stage 1/2. They live in 1Y3 rack and AUX_OMC_NORTH rack.

2. Restored all IFO optics alignment to the position where I aligned this afternoon (for SRM, I didn't aligned it; it restored at the saved value on May 26).

3. Centered all the oplevs. They can be used for a reference for alignment change before and after the vent.

I will leave PSL mechanical shutter and green shutters closed just in case.

Some MEDM screenshots below.
MEDMscreenshotswithCOW_20120629.png

ELOG V3.1.3-